# Augmented doubly robust post-imputation inference for proteomic data

### v1

@@ 3.2 Simulation study
We investigate the performance of our method compared to several other methods on simulated data. Six methods are compared; Full, Complete, MICE, DR_W, DR_UW (proposed), and Plugin, where they differ in the approach to obtain $P_{1},\ldots,P_{p}$ in Algorithm 1. The Full method uses the practically unavailable data $\mathbf{Y}$ without missingness. We perform a linear regression for each column of $\mathbf{Y}$ on low-dimensional covariates $\mathbf{W}=(a,x)$, where $a$ is the variable of interest and $x$ represents any other covariates. We then use a linear regression $t$-test to decide if the coefficient of $a$ equals zero. The Complete method works in the same way as the Full method, but uses only observed samples. The MICE method uses a multiple imputation method to impute missing values and then performs the test if the coefficient of $a$ is equal to zero using a statistic proposed by Rubin (1987). MICE is not computationally feasible when high-dimensional auxiliary variables are used for imputation. For this reason, missing values are imputed on the basis of low-dimensional covariates only. The DR_W method is similar to Algorithm 1, but the columns of $\hat{\bm{\nu}}$ are fitted by a linear regression model only with low-dimensional covariates. The DR_UW method follows Algorithm 1. The Plugin method regresses each column of the fitted outcomes $\hat{\bm{\nu}}$ in Algorithm 1 on the low-dimensional variables and performs a linear regression $t$-test for the coefficient of $a$. All five methods, except for the MICE method, require a choice of variance estimator to perform the linear regression $t$-test. For the Full, Complete, and Plugin methods, the usual OLS variance estimator is used. For the DR_W and DR_UW method, either the usual OLS variance estimator (in Models 1 and 2 below) or the heteroskedastic-consistent estimator (5) (in Models 3 and 4) is used. The six methods repeat the same procedure to obtain the p-values for each column of the outcome matrix. Then we transform the p-values into the Benjamini-Hochberg q-values and select the indices whose q-values are less than $\alpha=0.3$ to identify the discoveries. For each of the six methods, the fraction of false discoveries over the number of total discoveries (FDR; False Discovery Rate) and the fraction of true discoveries over the number of signal peptides (TPR; True Positive Rate) are reported. An ideal method would control FDR within $\alpha$, and have a TPR close to one. Two sample sizes of $n=200,500$ and a dimension $p=1000$ are considered. The number of repetitions is 200.
@@
    Gaussian data without $X$ (MCAR); $y^{j}_{i}=0.3s^{j}a_{i}+\epsilon^{j}_{i}$
@@
    Gaussian data (MCAR); $y^{j}_{i}=x_{i}+0.3s^{j}a_{i}+\epsilon^{j}_{i}$
@@
    Gaussian data (MAR); $y^{j}_{i}=x_{i}+0.3s^{j}a_{i}+\epsilon^{j}_{i}$
@@
    Skewed data (MAR); $y^{j}_{i}=x_{i}+0.08s^{j}a_{i}+\epsilon^{j}_{i}$
Correlation between peptides is simulated using a realistic covariance structure to model the noise terms associated with each peptide; The covariance ($\Sigma$) was estimated from peptides measured in brain tissue (MacDonald et al., 2017). For Models 1, 2 and 3, we simulate $n$ i.i.d. vectors $(\epsilon_{i}^{1},\ldots,\epsilon_{i}^{p})$, $i=1,\ldots,n$, using a multivariate normal distribution with zero mean and covariance $\Sigma$. For Model 4, we generate skewed noise as follows: simulate multivariate normal variables as before, for each peptide add a constant to ensure that all entries are positive, apply a log transformation, and finally recenter each peptide at zero. Covariates $x_{1},\cdots,x_{n}\in\mathbb{R}$ are generated independently from a uniform distribution in $(0,1)$. After generating covariates and noise, each outcome $y^{j}_{i}$ is randomly masked with the probability determined from the missingness model. In Models 1 and 2, each $y^{j}_{i}$ is missing completely at random (MCAR) with equal probability: $\mathbb{P}(C_{ij}=0)=0.3$. In Models 3 and 4, $y^{j}_{i}$ is missing at random (MAR): $\mathbb{P}(C_{ij}=0)={e^{x_{i}}}/\{2(1+e^{x_{i}})\}$.
Figure 2: Performance of different methods on simulated data according to Model 3.
Figure 2 summarizes the FDR and TPR of six methods applied to Model 3. As expected, FDR is well controlled for the Full, Complete, MICE, DR_W, and DR_UW methods, whereas the Plugin method inflates the FDR. This occurs because the differences between the case and control data in signal peptides bleed into correlated null peptides during the imputation procedure. When no signals are injected into any of the peptides or when the signal is carried for sets of correlated peptides, the Plugin method is also well-controlled (results not shown).
@@ 4 Single-cell protein abundance varies with cell size
We analyze a single-cell proteomic dataset based on mass spectrometry (Leduc et al., 2022) with the primary interest of detecting peptides whose abundance varies strongly with cell size. A peptide is a short chain of amino acids that constitute proteins and is a useful unit for quantitative analysis. The impact of cell size on cell physiology is of interest in two domains: large cell size may be a cause rather than a consequence of cell senescence (Lanz et al., 2022; Jones et al., 2023); and cell size is a determinant of stem cell fate (Lengefeld et al., 2021). The covariates for the analysis of these data are cell type (melanoma or monocyte) and three continuous variables (diameter, digestion, and elongation). These four variables form a low-dimensional covariate $\mathbf{W}$.
@@
The proportion of observed cells differs significantly between the peptides, and for 85.6% of the peptides, the observation rate is smaller than 0.5. When the observation rate is too low, it is not reasonable to expect that any method will perform satisfactorily; therefore, we focus on peptides with a missingness of no more than 50%. Because the threshold for the observation rate is controversial, for the main analysis, we provide a range of results with respect to different thresholds (0.5, 0.6, 0.7, 0.9). For other parts of the analysis, including exploratory data analysis and realistic simulation, we focus on a threshold of 0.7. After removing peptides whose observation rates are less than 0.7, there are a total of 753 remaining peptides.
@@
We first present exploratory data analysis to provide a rational basis for applying our method. The distribution of cell-wise peptide abundance data for peptides with more than 70% of observed rates reveals a Gaussian-like distribution (Fig. 3 A); thus, these data are well suited to our imputation model, which is a VAE model tailored to a Gaussian distribution; see Appendix B for more details. The distribution of pairwise distances between cells before and after imputation shows a noticeable reduction in distances after imputation, indicating that the overall variance of abundance data has decreased (Fig. 3B). Next, we check the assumption of MAR by examining the relationship between the measured variables and the observed cell-wise rate among the peptides (propensity score). The lack of a relationship between the residuals of the estimated propensity score and the mean abundance in cells, after regressing each of them into four covariates, shows that the observed relationship between the abundance of the peptide and the propensity score is largely explained by the measured covariates (Fig. 3C). However, the joint distribution of the cell-specific propensity score and each covariate, along with its marginal distribution, illustrates that each covariate is related to the propensity score to some extent, supporting an analysis under the assumption of MAR (Fig. 3 D). Furthermore, a reasonable imputation model can be built upon the robust relationship between peptides. Examining the quantile value of 0.9 of the absolute correlation coefficient for each peptide with other peptides reveals a strong correlation pattern. These values generally fall between 0.1 and 0.5, providing a good foundation for a high-dimensional imputation model (see Figure C4 in Appendix C for details).
@@
This section is organized as follows. First, we conduct realistic simulations to see how the proposed method works on this dataset with some artificially generated ground truth. Next, we present the results for the main analysis, where the primary focus of our analysis is to identify peptides whose abundance varies with the diameter of the cell. For all settings, we compare the proposed DR_UW method with the Complete, DR_W, and Plugin methods. The Full method is not considered due to a lack of data.
@@ 4.1 Realistic simulation
Before going into the main analysis, we check the performance of the proposed method with some artificially generated ground truth incorporated into this dataset and compare different methods in terms of TPR and FDR. Specifically, we artificially generate the type variable (case control) while keeping all other variables unchanged. In Setting 1, we randomly permute the measured type variable among the cells. In Setting 2, the type variable is randomly generated from a Bernoulli distribution with a probability proportional to the propensity score. Signal peptides are randomly selected for 10% of the total considered peptides, and then a positive signal generated from a normal distribution with mean 0.2 and variance 0.05 is added to the case cells in the signal peptides. After imputation, we identify peptides with different cutoffs for the q-value, 0.01, 0.05, and 0.3.
The Complete, DR_W, and DR_UW methods detect a reasonable proportion of signal peptides while maintaining control of FDR in both settings. For FDR=0.01 and 0.05, the DR_UW method provides better TPR than the Complete and DR_W method. The Complete method is better than the DR_W method in TPR, indicating that low-dimensional imputation is too noisy for satisfactory results. For FDR=0.3, all three methods, Complete, DR_W, and DR_UW, achieve near-perfect TPR. The Plugin method severely inflates FDR for all FDR levels, and the TPR is lower than the other three methods. These results are summarized in Figure C5 and C6 in Appendix C.
@@ 4.2 Main analysis
One objective of this analysis is to detect peptides whose abundance varies strongly with cell size. We first filter the peptides by applying varying thresholds (0.5, 0.7, 0.9) to the observation rates of the peptides and focus on analyzing those peptides. A larger proportion of peptides, whose observed rates are greater than 0.2, is used to feed the imputation procedure. After imputation, peptides are selected based on linear regression models:
@@
where we compute the p-values associated with the diameter variable. The p-values are transformed into q-values using the BH procedure. Based on estimated coefficients and the corresponding q-values, the Complete, DR_W, and DR_UW methods exhibit roughly similar distribution patterns; however, the Plugin method presents inflation of q-values compared to the other three methods due to the signal bleeding effect (Fig. 4).
@@
Figure 4: Volcano plot of peptide discoveries by different methods in the single-cell proteomics dataset analyzed in Section 4 when an observability threshold of 0.7 is applied. The red line indicates the q-value cutoff of 0.05.
Peptide discoveries vary according to the different methods and the different threshold settings. For the purpose of comparison, we derive a robustness metric under the assumption that meaningful relationships between peptide abundance and cell diameter are strictly positive. It follows that the estimated beta coefficients tend to be positive for the signal peptides and symmetrically distributed about zero for the null peptides. This assumption appears to be valid as there is a notably distinct pattern of positive and negative signs of the beta values (Figure 4). Several recent papers (Dai et al., 2023; Guo et al., 2023; Du et al., 2023) derive an FDR metric from a statistic that possesses such an asymmetric structure of null and non-null (signal) scenarios. Motivated by these, we calculate empirical FDR as the number of significant peptides with $\hat{\bm{\beta}}<0$ divided by the number of significant peptides with $\hat{\bm{\beta}}>0$. We find that this metric gives a reasonable interpretation throughout the various settings. An ideal discovery result has a lower value of empirical FDR and a larger number of discoveries.
@@
The performance of each method is evaluated in settings with different thresholds for observation rate and q-values (Table 1). We first fix the q-value cutoff at 0.05 and apply different thresholds to the observation rates of peptides. For a threshold of 0.9, empirical FDR is well-controlled for the Complete, DR_W and DR_UW methods, but the number of discoveries is relatively small because many peptides are excluded from the analysis. The Plugin method provides the largest discoveries, but its empirical FDR is inflated. When the threshold is lowered to 0.7, 0.6, and 0.5, the number of discoveries becomes larger, and empirical FDR tends to increase. This is natural because if the peptides with a high rate of missingness are introduced, the inference problem becomes more challenging. However, compared to the Plugin method, the other three methods consistently give better control of the empirical FDR, the proposed method DR_UW being the best. In addition, DR_UW provides a larger number of discoveries. This is consistent with what we observed from the simulations. It controls the FDR well while achieving greater TPR. The plugin method provides the largest number of discoveries, but it generally inflates the empirical FDR. Similar results hold when we fix the threshold to 0.7 and apply different q-value cutoffs.
@@
Observability threshold q-value cutoff Empirical FDR Number of selected peptides
@@
Table 1: Empirical FDR and number of peptides selected with each method under different combinations of thresholds applied to an observed rate of peptides and q-value cutoffs.
To further verify the robustness of the result, we map the peptides discovered by DR_W, DR_UW, and the Plugin method to the corresponding proteins, and check their overlaps with the discoveries of the Complete method (a q-value cutoff 0.05) at the protein level. We assess their contributions by only considering additional discoveries beyond those made by the Complete method with a q-value cutoff of 0.01. A threshold 0.7 is applied to the observation rate of peptides. When applying a q-value of 0.01, the additional discoveries of the DR_UW method are largely robust at the protein level; 90% of them overlap with those discovered by the Complete method. As we increase the q-value cutoff to 0.05, 0.1, and 0.3, the proportion of such overlaps tends to decrease, but the number of additional discoveries is much larger. Across all q-value cutoffs, the DR_UW method consistently provides more additional discoveries than the DR_W method, and has a slightly smaller proportion of overlaps with the Complete method. This aligns with theory and simulation results in the sense that both doubly robust methods exhibit reasonable control of the false discovery rate, while DR_UW demonstrates better efficiency. The Plugin method discovers the largest number of additional peptides, but as expected, its findings do not largely overlap with the discoveries of the Complete method at the protein level. The detailed results are summarized in Table C1 in Appendix C

### v2

@@ 3.2 Simulation study
We investigate the performance of our method compared to several other methods on simulated data. Eight methods are compared; Full, Complete, MICE, SVD, MissForest, DR_W, DR_UW (proposed), and Plugin, where they differ in the approach to obtain $P_{1},\ldots,P_{p}$ in Algorithm 1. The Full method uses the practically unavailable data $\mathbf{Y}$ without missingness. We perform a linear regression for each column of $\mathbf{Y}$ on low-dimensional covariates $\mathbf{W}=(a,x)$, where $a$ is the variable of interest and $x$ represents any other covariates. We then use a linear regression $t$-test to decide if the coefficient of $a$ equals zero. The Complete method works in the same way as the Full method, but uses only observed samples. The MICE method uses a multiple imputation method to impute missing values and then performs the test if the coefficient of $a$ is equal to zero using a statistic proposed by Rubin (1987).
MICE is not computationally feasible when high-dimensional auxiliary variables are used for imputation. For this reason, missing values are imputed on the basis of low-dimensional covariates only. The SVD and MissForest methods use an approach we call “Plugin-missing”, where only the missing values are imputed while the observed values remain unchanged. SVD estimates the missing values as a linear combination of the 10 most significant peptides (Troyanskaya et al., 2001), and MissForest makes predictions iteratively based on random forests until convergence (Stekhoven and Bühlmann, 2012). The DR_W method is similar to Algorithm 1, but the columns of $\hat{\bm{\nu}}$ are fitted by a linear regression model only with low-dimensional covariates. The DR_UW method follows Algorithm 1. The Plugin method regresses each column of the fitted outcomes $\hat{\bm{\nu}}$ in Algorithm 1 on the low-dimensional variables and performs a linear regression $t$-test for the coefficient of $a$. All seven methods, except for the MICE method, require a choice of variance estimator to perform the linear regression $t$-test. For the Full, Complete, Plugin-missing and Plugin methods, the usual OLS variance estimator is used. For the DR_W and DR_UW method, either the usual OLS variance estimator (in Models 1 and 2 below) or the heteroskedastic-consistent estimator (5) (in Models 3 and 4) is used. The eight methods repeat the same procedure to obtain the p-values for each column of the outcome matrix. Then we transform the p-values into the Benjamini-Hochberg q-values and select the indices whose q-values are less than a predefined cutoff $\alpha$ to identify the discoveries. For each of the eight methods, the fraction of false discoveries over the number of total discoveries (FDR; False Discovery Rate) and the fraction of true discoveries over the number of signal peptides (TPR; True Positive Rate) are reported. An ideal method would control FDR within $\alpha$, and have a TPR close to one. Two sample sizes of $n=200,500$ and a dimension $p=1000$ are considered. The number of repetitions is 200.
Remark 3.1.
The Plugin-missing approach lies between the Plugin and the Complete methods and is another widely used technique in practice, especially when the imputation algorithm only provides an estimate for the missing values. For the VAE result, we only consider a combination with the Plugin method because it is the standard approach when the imputation algorithm outputs estimates for the entire matrix. This is sometimes referred to as a denoising procedure; Wang et al. (2019) provides a comparative analysis of these procedures.
@@
    Gaussian data without $X$ (MCAR); $y^{j}_{i}=c_{1}s^{j}a_{i}+\epsilon^{j}_{i}$
@@
    Gaussian data (MCAR); $y^{j}_{i}=x_{i}+c_{1}s^{j}a_{i}+\epsilon^{j}_{i}$
@@
    Gaussian data (MAR); $y^{j}_{i}=x_{i}+c_{1}s^{j}a_{i}+\epsilon^{j}_{i}$
@@
    Skewed data (MAR); $y^{j}_{i}=x_{i}+c_{2}s^{j}a_{i}+\epsilon^{j}_{i}$
Correlation between peptides is simulated using a realistic covariance structure to model the noise terms associated with each peptide; The covariance ($\Sigma$) was estimated from peptides measured in bulk brain tissue (MacDonald et al., 2017). For Models 1, 2 and 3, we simulate $n$ i.i.d. vectors $(\epsilon_{i}^{1},\ldots,\epsilon_{i}^{p})$, $i=1,\ldots,n$, using a multivariate normal distribution with zero mean and covariance $\Sigma$. For Model 4, we generate skewed noise as follows: simulate multivariate normal variables as before, for each peptide add a constant to ensure that all entries are positive, apply a log transformation, and finally recenter each peptide at zero. Covariates $x_{1},\cdots,x_{n}\in\mathbb{R}$ are generated independently from a uniform distribution in $(0,1)$. After generating covariates and noise, each outcome $y^{j}_{i}$ is randomly masked with the probability determined from the missingness model. In Models 1 and 2, each $y^{j}_{i}$ is missing completely at random (MCAR) with equal probability: $\mathbb{P}(C_{ij}=0)=0.3$. In Models 3 and 4, $y^{j}_{i}$ is missing at random (MAR): $\mathbb{P}(C_{ij}=0)={e^{x_{i}}}/\{2(1+e^{x_{i}})\}$. We use different signal strengths for different sample sizes to construct meaningful comparisons between methods; $c_{1}=0.4$, $c_{2}=0.12$ for $n=200$ and $c_{1}=0.3$, $c_{2}=0.08$ for $n=500$.
Figure 2: Performance of different methods on simulated data according to Model 3. The X-axis represents different levels of q-value cutoffs, and the Y-axis represents either FDR or TPR.
Figure 2 summarizes the FDR and TPR of eight methods applied to Model 3. As expected, FDR is well controlled for the Full, Complete, MICE, DR_W, and DR_UW methods, whereas the Plugin method inflates the FDR. This occurs because the differences between the case and control data in signal peptides bleed into correlated null peptides during the imputation procedure. When no signals are injected into any of the peptides or when the signal is carried for sets of correlated peptides, the Plugin method is also well-controlled (results not shown).
@@
Two Plugin-missing methods control the FDR, but the TPR is even lower than that of the Complete method, indicating that there is no benefits from the additional simulation step. When stronger signals are injected (i.e., $y^{j}_{i}=x_{i}+s^{j}a_{i}+\epsilon^{j}_{i}$ in Model 3), they have inflated FDR (Figure C4). Thus, the Plugin-missing method performs somewhat between the Plugin method and the Complete method, which is not surprising, as the response in the Plugin-missing methods only uses a partial substitution of the original data.
@@ 4 Single-cell protein abundance varies with cell size
For illustration, we analyze a single-cell proteomic dataset based on mass spectrometry as published by Leduc et al. (2022). Data processing, as described by the authors, involves several steps, including QC and a log transformation. The primary interest is detecting peptides whose abundance varies strongly with cell size. A peptide is a short chain of amino acids that constitute proteins and is a useful unit for quantitative analysis. The proportion of observed cells differs significantly between the peptides, and for 85.6% of the peptides, the observation rate is smaller than 0.5. When the observation rate is too low, it is not reasonable to expect that any method will perform satisfactorily; therefore, we focus on peptides with a missingness of no more than 50%. Because the threshold for the observation rate is controversial, for the main analysis, we provide a range of results with respect to different thresholds (0.5, 0.6, 0.7, 0.9). For other parts of the analysis, including exploratory data analysis and realistic simulation, we focus on a threshold of 0.7. After removing peptides whose observation rates are less than 0.7, there are a total of 753 remaining peptides.
@@
We first present exploratory data analysis to provide a rational basis for applying our method. The distribution of cell-wise peptide abundance data for peptides with more than 70% of observed rates reveals a Gaussian-like distribution (Fig. 3A); thus, these data are well suited to our imputation model, which is a VAE model tailored to a Gaussian distribution; see Appendix B for more details. The distribution of pairwise distances between cells before and after imputation shows a noticeable reduction in distances after imputation, indicating that the overall variance of abundance data has decreased (Fig. 3B). Next, we check the assumption of MAR by examining the relationship between the measured variables and the observed cell-wise rate among the peptides (propensity score). Four measured variables are examined: cell diameter, elongation, type, and digest. The lack of a relationship between the residuals of the estimated propensity score and the mean abundance in cells, after regressing each of them into four covariates, shows that the observed relationship between the abundance of the peptide and the propensity score is largely explained by the measured covariates (Fig. 3C). However, the joint distribution of the cell-specific propensity score and each covariate, along with its marginal distribution, illustrates that each covariate is related to the propensity score to some extent, supporting an analysis under the assumption of MAR (Fig. 3D). Furthermore, a reasonable imputation model can be built upon the robust relationship between peptides. Examining the quantile value of 0.9 of the absolute correlation coefficient for each peptide with other peptides reveals a strong correlation pattern. These values generally fall between 0.1 and 0.5, providing a good foundation for a high-dimensional imputation model (see Figure C6 for details).
@@
This section is organized as follows. First, we conduct realistic simulations to see how the proposed method works on this dataset with an artificially generated ground truth. Next, we present the results for the main analysis, where the primary focus of our analysis is to identify peptides whose abundance varies with the diameter of the cell. For all settings, we compare the proposed DR_UW method with the Complete, DR_W, and Plugin methods. Naturally, the Full method cannot be considered.
@@ 4.1 Realistic simulation
Before going into the main analysis, we check the performance of the proposed method with some artificially generated ground truth incorporated into this dataset and compare different methods in terms of TPR and FDR. Specifically, we artificially generate the type variable (case and control) while keeping all other variables (cell diameter, elongation and digest) unchanged. In Setting 1, we randomly permute the measured type variable among the cells. In Setting 2, the type variable is randomly generated from a Bernoulli distribution with a probability proportional to the propensity score. Signal peptides are randomly selected for 10% of the total considered peptides, and then a positive signal generated from a normal distribution with mean 0.2 and variance 0.05 is added to the case cells in the signal peptides. After imputation, we identify differentially abundant peptides using various q-value cutoffs of 0.01, 0.05, and 0.3.
The Complete, DR_W, and DR_UW methods detect a reasonable proportion of signal peptides while maintaining control of FDR in both settings. For FDR=0.01 and 0.05, the DR_UW method provides better TPR than the Complete and DR_W method. The Complete method is better than the DR_W method in TPR, indicating that low-dimensional imputation is too noisy for satisfactory results. For FDR=0.3, all three methods, Complete, DR_W, and DR_UW, achieve near-perfect TPR. The Plugin method severely inflates FDR for all FDR levels, and the TPR is lower than the other three methods. These results are summarized in Figure C7 and C8 in Appendix. Notably, the improvement in results is not as striking as those displayed in Figure 2, which are based on bulk tissue simulations. The value of the proposed method depends on the quality of imputation we can obtain from a dataset. As single-cell data are more difficult to impute due to a higher noise level and greater missingness, it is not surprising that the performance is somewhat diminished.
@@ 4.2 Main analysis
One objective of this analysis is to detect peptides whose abundance varies strongly with cell size. We first filter the peptides by applying varying thresholds (0.5, 0.6, 0.7, 0.9) to the observation rates of the peptides and focus on analyzing those peptides. A larger proportion of peptides, whose observed rates are greater than 0.2, is used to feed the imputation procedure. After imputation, peptides are selected based on linear regression models:
@@
where we compute the p-values associated with the diameter variable. The p-values are transformed into q-values using the BH procedure. Based on estimated coefficients and the corresponding q-values, the Complete, DR_W, and DR_UW methods exhibit roughly similar distribution patterns; however, the Plugin method yields a larger number of significant q-values compared to the other three methods due to the signal bleeding effect (Fig. 4).
@@
Figure 4: Volcano plot of peptide discoveries by different methods in the single-cell proteomics dataset analyzed in Section 4 when an observability threshold of 0.9 is applied. The horizontal line indicates the q-value cutoff of 0.05.
The objective of this analysis is to detect peptides whose abundance varies strongly with cell size. The impact of cell size on cell physiology is of interest in two domains: large cell size may be a cause rather than a consequence of cell senescence (Lanz et al., 2022; Jones et al., 2023); and cell size is a determinant of stem cell fate (Lengefeld et al., 2021). Cells develop various shapes and sizes to perform their specific functions and the coordination of cell growth and division to achieve and maintain these cellular phenotypes has been a key research interest in cell biology (Turner, Ewald and Skotheim, 2012; Lloyd, 2013; Liu, Yan and Kirschner, 2024). Because large cells need to produce and maintain more cellular components than small cells, it is anticipated that cell size will be positively associated with some protein abundances, and this is numerically validated in multiple earlier studies (Marguerat and Bähler, 2012; Lanz et al., 2022; Jones et al., 2023); however, there could be exceptions. For instance, a small cell that is growing rapidly might produce more of the proteins that are essential for this current developmental phase, and this phenomenon would lead to a small number of negative associations. Regardless of the direction, identifying the proteins associated with cell size could provide insights into abnormal cell growth, such as in cancer or stem cell fate (Lengefeld et al., 2021; Jones et al., 2023).
If the majority of true signals are positive, it follows that the estimated $\beta$ coefficients will tend to be positive for the signal peptides and symmetrically distributed around zero for the null peptides. Therefore, most of the significant coefficients will be positive, and only a small number of noisy null peptides will contribute to both positive and negative signs of the selected coefficients. We propose the mirror rate, defined as the number of significant peptides with $\hat{\bm{\beta}}<0$ divided by the number of significant peptides with $\hat{\bm{\beta}}>0$, as a metric to compare the reliability of the findings. As more noisy null peptides are included, the mirror rate will increase. Several recent papers (Dai et al., 2023; Guo et al., 2023; Du et al., 2023) derive an FDR metric from a statistic that exhibits such an asymmetric structure of null and non-null (signal) scenarios. In practice, we anticipate that our mirror rate may overestimate the false discovery rate due to the possibility of some proteins having a negative relationship. However, since such a fraction of relationships are commonly present in all the methods we compare, we expect it to provide a useful comparison across methods.
@@
The performance of each method is evaluated in settings with different thresholds for observation rate and q-values (Table 1). We first fix the q-value cutoff at 0.05 and apply different thresholds to the observation rates of peptides. For a threshold of 0.9, mirror rates are well-controlled for the Complete, DR_W and DR_UW methods, but the number of discoveries is relatively small because many peptides are excluded from the analysis. The Plugin method provides the largest discoveries, but its mirror rate is inflated. When the threshold is lowered to 0.7, 0.6, and 0.5, the number of discoveries becomes larger, and mirror rate tends to increase. This is natural because if the peptides with a high rate of missingness are introduced, the inference problem becomes more challenging. However, compared to the Plugin method, the other three methods consistently give better control of the mirror rate, the proposed method DR_UW being the best. In addition, DR_UW provides a larger number of discoveries. This is consistent with what we observed from the simulations. It controls the FDR well while achieving greater TPR. The Plugin method provides the largest number of discoveries, but it generally inflates the mirror rate. Similar results hold when we fix the threshold to 0.7 and apply different q-value cutoffs.
@@
Observability threshold q-value cutoff Mirror rate Number of selected peptides
@@
Table 1: Mirror rate and number of peptides selected with each method under different combinations of thresholds applied to an observed rate of peptides and q-value cutoffs.
To further verify the robustness of the result, we map the peptides discovered by DR_W, DR_UW, and the Plugin method to the corresponding proteins, and check their overlaps with the discoveries of the Complete method (a q-value cutoff 0.05) at the protein level. We assess their contributions by only considering additional discoveries beyond those made by the Complete method with a q-value cutoff of 0.01. A threshold of 0.7 is applied to the observation rate of peptides. When applying a q-value of 0.01, the additional discoveries of the DR_UW method are largely robust at the protein level; 90% of them overlap with those discovered by the Complete method. As we increase the q-value cutoff to 0.05, 0.1, and 0.3, the proportion of such overlaps tends to decrease, but the number of additional discoveries is much larger. Across all q-value cutoffs, the DR_UW method consistently provides more additional discoveries than the DR_W method, and has a slightly smaller proportion of overlaps with the Complete method. This aligns with theory and simulation results in the sense that both doubly robust methods exhibit reasonable control of the false discovery rate, while DR_UW demonstrates better efficiency. The Plugin method discovers the largest number of additional peptides, but as expected, its findings do not largely overlap with the discoveries of the Complete method at the protein level. The detailed results are summarized in Table C1.